Opened 9 years ago

Closed 7 years ago

#356 closed flight processing (fixed)

EUFAR10/03 (A.NEW), flight day 196/2010,Lisbon

Reported by: ella Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by iopa)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_03-2010_196_Lisbon

Data arrived from ARSF via network transfer on 2nd/08/2010.

Scientific objectives:
Objective 1.- Development of a reliable tool to estimate coastal water quality indicators by using remote sensing
techniques, both airborne and satellite.
1.1. | Development of a set of semi-empirical algorithms based on hyper-spectral airborne and in-situ data for
estimating the concentrations of the water constituents;
1.2. | Migration of the developed algorithms into the satellite workflow.

Priority: a4h

PI: Jose da Silva

Sensors:

  • Eagle (22/12/2010)
  • Hawk (22/12/2010)
  • Leica LIDAR (03/12/2010)
  • RCD Camera (03/12/2010)

Attachments (2)

line_06.png (51.8 KB) - added by emca 8 years ago.
line 06
1962010.sync (236 bytes) - added by iopa 8 years ago.
Eagle & SCT offsets

Download all attachments as: .zip

Change History (45)

comment:1 Changed 9 years ago by ella

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:2 Changed 9 years ago by ella

Starting navigation processing.

comment:3 Changed 8 years ago by jb362

IPAS navigation processing done.

comment:4 Changed 8 years ago by ella

Waiting for logsheet.

Applanix processing was problematic because the flight was 100km away from the nearest base station.
POSproc failed to create an Sbet file.

comment:5 Changed 8 years ago by jb362

Starting camera processing.

comment:6 Changed 8 years ago by emca

Starting LiDAR processing

Changed 8 years ago by emca

line 06

comment:7 Changed 8 years ago by emca

There are six LiDAR lines for this project.

Lines 1 and 2 are very short, contain very little data (approx 30% first returns) and are very noisy. From the leica flightline log it it looks like there was also an issue with the range correction during the capture of these lines.

Lines 3 and 4 contain no useful data.

Line 5 is partially over water and contains a 90 degree turn.

Line 6 consists of four distinct lines, including the turns in between (see attached screenshot). This may need to be separated into 4 lines.

There is little overlap between lines (just the cross of Line 6 with itself) which will make correcting for roll difficult.

Going to wait until next week when ARSF Ops return and we can get the logsheet before proceeding any further with this project.

comment:8 Changed 8 years ago by jb362

Started camera processing for this project but in the kml file it's very hard to see which images will be needed and which won't. This is because it the camera wasn't turned off between flight lines. This processing will therefore be held until we have the logsheet for this project. If I have left before we do get the logsheet iopa will be picking up the camera processing.

comment:9 Changed 8 years ago by jb362

Also, all the .tif files for the camera processing have already been created.

comment:10 Changed 8 years ago by emca

Call from Phil to say there are still looking for the logs sheets for this project as well as for Lisbon flight day 197

Both projects were flown on an ad hoc basis. They were unable to turn the LiDAR off during the flight as they were flying over water but the hyperspectral was switched off between lines to conserve disk space. As a result there will not be a match between the eagle/hawk and the LiDAR lines.

comment:11 Changed 8 years ago by emca

Logsheet has arrived and been saved in ~arsf/arsf_data/2010/flight_data/archive-overflow/EUFAR10_03-2010_196_Lisbon/admin

comment:12 Changed 8 years ago by jb362

Iopa will be taking over this camera project.

comment:13 Changed 8 years ago by ella

Starting camera delivery check.

comment:14 Changed 8 years ago by ella

Camera check complete.
Two photos have been removed and the read_me and kml file have been updated.
Ready to deliver.

comment:15 Changed 8 years ago by iopa

  • Description modified (diff)

RCD camera was dispatched today, 16th September 2010.

comment:16 Changed 8 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking over hyper-spectral processing for this project.

comment:17 Changed 8 years ago by knpa

Issue with applanix - raw data in two parts with a gap in the middle. This gap covers one of the lines (Gen3).

Also, cannot find eagle data for first two lines (Gen 1/2) and (Gen 3).

comment:18 Changed 8 years ago by knpa

  • Description modified (diff)

Move to repository, can be rsynced.

comment:19 Changed 8 years ago by knpa

Phone call from Phil to Ben:
The data collection for this project starts at 11:39 and earlier lines are tests.

That excludes Gen1/2 and Gen3, meaning that this project now has 5 lines containing eagle and hawk data that will be processed.

comment:20 Changed 8 years ago by knpa

SCTs impossible to correct on this project as almost all lines are over water.
H6 had no sync but managed to correct gross sync issues by matching water patterns with eagle.

SCTs

Eagle Hawk
Line 1 0 0
Line 2 0 0
Line 3 0 0
Line 4 0 0
Line 5 0 1.5 (sync)

comment:21 Changed 8 years ago by knpa

Correction: above Hawk 5 value should be 0.15 NOT 1.5.

comment:22 Changed 8 years ago by knpa

First 12 bands of Hawk lev1s are blank. The raws are fine. Appears to be an issue with the bandset labels in the header files, they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:23 Changed 8 years ago by benj

  • Description modified (diff)

Corrected PI (recorded wrong on data unpacking)

comment:24 Changed 8 years ago by knpa

Eagle processing has been completed.

PI has said they want to wait until until both hyperspectral datasets can be delivered together (i.e. until hawk is ready in December at earliest).
(e-mail 25/10/10).

comment:25 Changed 8 years ago by emca

LiDAR data ready for delivery check at
~airborne/workspace/EUFAR10_03-2010_196_Lisbon/delivery/20101109/EUFAR10_0

LiDAR data for lines for UL005 & UL0066 only. UL006 was very long so it has been split into three to correspond roughly to the three Eagle/Hawk lines.

Note: It was not possible to correct for roll on this project. Did not create a DEM from the LiDAR data as it is mostly over water.

comment:26 Changed 8 years ago by benj

  • Description modified (diff)

comment:27 Changed 8 years ago by iopa

LiDAR has been checked,

the data and directories are fine considering there is no DEM.
there is only a typo in the readme in the data quality section where it should be 'manageable'

otherwise LiDAR is ready.

comment:28 Changed 8 years ago by emca

  • Description modified (diff)

LiDAR and Photography dispatched 03/12/2010

(note: photography is a re-delivery as data sent on 16/09/10 was sent to the wrong PI)

comment:29 Changed 8 years ago by knpa

iopa reprocessing hawk

comment:30 Changed 8 years ago by knpa

  • Owner changed from knpa to iopa

Some lines in this project are too large for tiffs to be created, need to give option to az to produce level 3 bil files instead.

comment:31 Changed 8 years ago by iopa

Processed Eagle and Hawk, delivery created, ready for checking.

comment:32 Changed 8 years ago by knpa

Taking hyperspectral delivery check.

comment:33 Changed 8 years ago by knpa

Delivery check complete. Need to address the following:

  • Need to rename flightlines on logsheet to match Readme (save original)
  • should mention that we could not correct for SCTs due to water (in non-technical language)
  • hawk blue banding in screenshot looks interesting - is that anything to worry about?
  • banded underflows for all hawk lines - coincides with screenshot?
  • overflows from sun-glint - eagle lines 1,2,3,4 (first ~40 bands)
  • overflows caused by land - hawk line 2 (and a bit in eagle)
  • large difference between actual size and header file size for e19601c
  • some lev1s not readable by all (should use chmod 644 apparently)

comment:34 Changed 8 years ago by iopa

Delivery corrected and data dispatched on 22/12/2010.

comment:35 Changed 8 years ago by iopa

  • Description modified (diff)

Changed 8 years ago by iopa

Eagle & SCT offsets

comment:36 Changed 8 years ago by anch

  • Summary changed from EUFAR10/03, flight day 196/2010,Lisbon to EUFAR10/03 (A.NEW), flight day 196/2010,Lisbon

comment:37 Changed 8 years ago by knpa

  • Owner changed from iopa to knpa

r-synced project.

Beginning archival.

comment:38 Changed 8 years ago by knpa

  • Component changed from Processing: general to Archiving

comment:39 Changed 8 years ago by knpa

tabi data present

comment:40 Changed 8 years ago by knpa

Tarball created. Made available to NEODC.

comment:41 Changed 8 years ago by knpa

Backed up by NEODC. Moved to archive disk.

comment:42 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion

comment:43 Changed 7 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

Note: See TracTickets for help on using tickets.